OpenStudy (anonymous):

A distribution center receives shipments from three different factories in quantities of 50, 35, and 20. Three times a product is selected at random. Fing the probability that none of the products came from the second factory

OpenStudy (turingtest):

each selection there is a probability of\[\frac{35}{50+35+20}=\frac{35}{105}=\frac13\]that the product will come from the second factory that means that there is a\[\frac23\]chance it did not after three trials the probability that none will come from the second factory is therefor\[\frac23\cdot\frac23\cdot\frac23=\frac8{27}\]

OpenStudy (turingtest):

welcome to open study :D